[Hallc_running] [New Logentry] Hall C RC update: Tuesday, March 12

gaskelld at jlab.org gaskelld at jlab.org
Tue Mar 12 11:15:01 EDT 2019

Logentry Text:
Yesterday, we completed all data-taking for the x=0.65, Q2=5.5 settings and moved on to the x=0.5, Q2=5.5 (pi-) setting yesterday evening.  This setting will likely be completed sometime swing shift today and we will move on to the equivalent pi+ setting.

Tomorrow morning, there will be an arc energy measurement (planned to start 8:30 am).

Beam delivery has continued to very efficient. - there were a couple interruptions yesterday due to issues with the power supply for the Hall C arc dipoles, but that seems to have been resolved since swing shift.

There will be no RC meeting today - status and run plan information can be found here:



This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3666746
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20190312/004b4384/attachment.html>

More information about the Hallc_running mailing list